DETAILED ANALYSIS OF THE 0^{+} − 0^{−} INVERSION DOUBLET IN D_{2}NCN and HDNCN CYANAMIDE

\underline{Z. KISIEL}, A. KRASNICKI, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; B. P. WINNEWISSER, M. WINNEWISSER, Department of Physics, The Ohio State University, Columbus, OH 43210.

The cyanamide molecule is one of the prototype systems for the study of the large-amplitude inversion motion at the nitrogen atom and it is also a known astrophysical species. The rotational and vibration-rotation spectrum of the parent species has already been studied in considerable detail.^{a}^{,}^{b} Information on the deuterated isotopologues has, hitherto, been limited to transitions measured at frequencies of less, or much less, than 60 GHz.

Presently we report a detailed analysis of pure rotational and vibration-rotation transitions in the 0^{+} − 0^{−} inversion doublets for the two deuterated species. The analysis is based on a 118-650 GHz spectrum for D_{2}NCN and HDNCN recorded at rotational resolution, and an 8-360 cm^{−1} spectrum of D_{2}NCN recorded with the Bruker IFS 120HR interferometer at resolution of 0.0025 cm^{−1}. The two spectra were assigned and analysed with the modified version of the AABS package^{c} for Assignment and Analysis of Broadband Spectra and the more extensive data set for D_{2}NCN contains well over 2000 lines with J ≤ 80 and K_{a} ≤ 15. The measured transition frequencies were fitted to within experimental accuracy with a unified scheme for fitting the inversion doublet transitions in H_{2}NCN, NHDCN and D_{2}NCN. The spectroscopic constants and information on the inversion potential determined from the spectra of the three isotopologues are compared and discussed.

aM. Birk et al., J. Mol. Spectrosc., 159, 69-78 (1993).

bG. Moruzzi et al. J. Mol. Spectrosc., 190, 353-364 (1998).

cZ. Kisiel et al., J. Mol. Spectrosc., 246, 39-56 (2007).